# Mirror-3DGS: Incorporating Mirror Reflections into 3D Gaussian Splatting

### v1

@@ Abstract
3D Gaussian Splatting (3DGS) has marked a significant breakthrough in the realm of 3D scene reconstruction and novel view synthesis. However, 3DGS, much like its predecessor Neural Radiance Fields (NeRF), struggles to accurately model physical reflections, particularly in mirrors that are ubiquitous in real-world scenes. This oversight mistakenly perceives reflections as separate entities that physically exist, resulting in inaccurate reconstructions and inconsistent reflective properties across varied viewpoints. To address this pivotal challenge, we introduce Mirror-3DGS, an innovative rendering framework devised to master the intricacies of mirror geometries and reflections, paving the way for the generation of realistically depicted mirror reflections. By ingeniously incorporating mirror attributes into the 3DGS and leveraging the principle of plane mirror imaging, Mirror-3DGS crafts a mirrored viewpoint to observe from behind the mirror, enriching the realism of scene renderings. Extensive assessments, spanning both synthetic and real-world scenes, showcase our method’s ability to render novel views with enhanced fidelity in real-time, surpassing the state-of-the-art Mirror-NeRF specifically within the challenging mirror regions. Our code will be made publicly available for reproducible research.
Keywords:
Gaussian Splatting Mirror Scene Novel View Synthesis
footnotetext: $\dagger$ means equal contribution.
1 Introduction
@@
Traditional methods [32, 36, 53, 31, 25] have attempted to simulate surface reflections by decomposing the material properties of objects and lighting conditions, thereby handling scenes with mirrors. However, these methods often face limitations in accurately inferring the surfaces and mirror components, struggling to effectively manage pure mirror reflections. With the advancement of Neural Radiance Field (NeRF) , some new methods [36, 14, 49] have been proposed. NeRFReN [14] enhances the rendering quality of scenes containing mirrors by employing independent radiance fields to simulate reflective and transmissive parts separately. Mirror-NeRF [49] utilizes Whitted Ray Tracing [41] to achieve physically accurate rendering of mirror reflections, achieving exceptional realism in novel view synthesis. However, the main challenges faced by these NeRF-based methods include long training times and slow rendering speeds, limiting their application in real-time rendering scenarios.
@@
Figure 1: The synthesized novel views and depth maps. 3DGS confuses mirror reflection, resulting in erroneous mirror depth. Ours properly handles content in the mirror with depth map almost identical to the ground truth (GT).
@@
Recently, 3D Gaussian Splatting (3DGS) [18] has emerged as a promising approach for real-time, high-fidelity 3D scene rendering. Instead of the slow volumetric rendering of NeRF, 3DGS explicitly represents scenes using 3D Gaussians and rasterizes 3D Gaussians onto the image plane using splatting techniques, achieving significant improvements in training and rendering speeds. However, 3DGS tends to treat mirror reflections as separate scenes, constructing Gaussian points for them as if they physically exist, leading to inaccurate capture of these virtual images within mirrors during reconstruction and rendering, affecting the realism and visual coherence of the scene, as shown in Fig. 1.
@@
To address these challenges, we propose Mirror-3DGS, a novel rendering framework based on 3DGS, aimed at producing high-fidelity novel views in mirror-containing scenes. Our Mirror-3DGS is centered on the principle that the virtual image presented on the mirror is equivalent to observing the real world from behind the mirror. In general, we first identify 3D Gaussians representing the mirror. Then, we observe from both the current viewpoint and its mirrored viewpoint, integrating views from these perspectives to get the final rendered image. To achieve this, we introduce an attribute to signify mirror-like properties to 3D Gaussians, and estimate the plane equation of the mirror based on the Gaussians with pronounced mirror properties, which allows us to derive the mirror transformation matrix and mirrored camera parameters. Our method involves a two-stage training process: 1) During the first stage, we learn a rough 3D Gaussian representation of the scene without content in the mirror and the mirror plane equation. We replace content within the mirror with a fixed color based on the ground truth mirror mask to better learn mirror attributes and prevent interference from virtual images. 2) In the second stage, we filter Gaussians with high mirror properties for precise estimation of the mirror equation. By integrating views from current and its mirrored viewpoints, we further optimize the rendering quality of the entire scene. To ensure the stability of the rendering process and improve the accuracy of plane equation estimation, we introduce planar consistency constraints and depth constraints through the whole training process. These constraints help achieving more realistic and refined rendering results. Experiments conducted across three synthetic and three real-world scenes demonstrate that our proposed Mirror-3DGS is capable of delivering rendering quality comparable to the state-of-the-art Mirror-NeRF, while significantly reducing training time and achieving real-time rendering. Our contributions can be summarized as follows:
@@
  • Innovative Rendering Framework for Mirror Scenes. We present Mirror-3DGS, an novel rendering framework that effectively addresses the inherent limitation of 3DGS in handling scenes with mirror, empowering it to achieve precise and realistic synthesis of novel views.
@@
  • Extensive Experiments. We conducted comprehensive experiments across both synthetic and real scenes to validate our proposed method. The results demonstrate our superiority in real-time rendering of high quality novel views, particularly surpassing the state-of-the-arts in mirror regions.
2 Related Work
2.1 Novel View Synthesis
The Neural Radiance Fields (NeRF) [27] significantly advanced the field of novel view synthesis, enhancing the capabilities of multi-view reconstruction. NeRF catalyzed an exponential surge in subsequent research efforts, focusing on pushing the frontiers of scene reconstruction across diverse domains. These endeavors have been geared towards refining reconstruction quality [2, 3, 38, 4, 15, 43], enhancing computational efficiency [37, 28, 33, 13, 30], enabling advanced editing functionalities [45, 23, 1, 40, 44], and progressing dynamic scene representation [29, 12, 8, 22, 10]. As a potent successor, 3D Gaussian Splatting (3DGS) [18] reconstructs and represents 3D scenes through the utilization of millions of 3D Gaussians. In contrast to explicit scene representations such as meshes and voxels, 3DGS excels in capturing complex geometries with an elevated level of rendering fidelity. Furthermore, when compared to prior implicit representations like NeRF, 3DGS presents unmatched advantages in terms of both training and rendering efficiency, thereby showcasing its utility in a variety of applications ranging from SLAM [17, 26] to 4D reconstruction [21, 24, 42, 46] and 3D content generation [34, 47].
2.2 Reflection-aware Rendering
Extensive research initiatives [5, 6, 7, 51, 52] have been dedicated to surmounting the complexities of accurately rendering physical reflections within 3D environments. Notably, RefNeRF [36] splits light into diffuse and specular components, employing a reflection-dependent radiance field for modeling reflections based on the view direction. SpecNeRF [25] introduces a learnable Gaussian directional encoding, enhancing the depiction of view-dependent effects under proximate lighting conditions. NeRFReN [14] segregates reflections from transmissions through distinct radiance fields, markedly improving renderings of scenes inclusive of mirrors. Both TraM-NeRF [35] and Mirror-NeRF [49] meticulously simulate mirror geometries and employ path tracing for rendering, delivering exceptional quality. However, the lingering issue of slow rendering speeds significantly hampers the broader applicability of these NeRF-based methods. The method proposed in this paper, grounded in the principles of plane mirror imaging, innovatively applies to modeling mirror scenes within the 3DGS framework. This strategy not only maintains 3DGS’s fast training and real-time rendering benefits but also adeptly navigates the challenges previously outlined, presenting a robust solution to mirror scene representation.
3 Methodology
3.1 Preliminaries
3DGS is an innovative and state-of-the-art approach in the field of novel view synthesis. Distinguished from implicit representation methods such as NeRF [27], which utilize volume rendering, 3DGS leverages the splatting technique [48] to generate images, achieving remarkable real-time rendering speed. Specifically, 3DGS represents the scene through a set of anisotropic Gaussians, defined with its center position $\boldsymbol{\mu}\in\mathbb{R}^{3}$, covariance $\mathbf{\Sigma}\in\mathbb{R}^{3\times 3}$ which can be decomposed into scaling factor $\boldsymbol{s}\in\mathbb{R}^{3}$ and rotation factor $\boldsymbol{r}\in\mathbb{R}^{4}$, color defined by spherical harmonic (SH) coefficients $\boldsymbol{c}\in\mathbb{R}^{3\times(k+1)^{2}}$ (where $k$ represents the order of spherical harmonics), and opacity $\alpha\in\mathbb{R}^{1}$. And the 3D Gaussian can be queried as follows:
$G(\mathbf{x})=e^{-\frac{1}{2}(\boldsymbol{x}-\boldsymbol{\mu})^{\top}\mathbf{\Sigma}^{-1}(\boldsymbol{x}-\boldsymbol{\mu})},$ (1)
@@ 3.1 Preliminaries
$\boldsymbol{\mu}^{\prime}={\mathbf{P}\mathbf{W}\boldsymbol{\mu}},$ (2)
$\mathbf{{\Sigma}^{\prime}}=\mathbf{J}\mathbf{W}\mathbf{\Sigma}\mathbf{W^{\top}}\mathbf{J^{\top}},$ (3)
where $\boldsymbol{\mu}^{\prime}$ and $\mathbf{{\Sigma}^{\prime}}$ separately represent the 2D mean position and covariance of the projected 3D Gaussian. $\mathbf{P}$, $\mathbf{W}$ and $\mathbf{J}$ donate the projective transformation, viewing transformation, and Jacobian of the affine approximation of $\mathbf{P}$, respectively. The color of the pixel on the image plane, denoted by $\mathbf{p}$, uses a typical neural point-based rendering [19, 20]. Let $\mathbf{C}\in\mathbb{R}^{H\times W\times 3}$ represent the color of rendered image where $H$ and $W$ represents the height and width of images, the rendering process outlined as follows:
${\mathbf{C}(\mathbf{p})}={\sum_{i=1}^{N}c_{i}\sigma_{i}\prod_{j=1}^{i-1}(1-\sigma_{j})},\quad\sigma_{i}=\alpha_{i}e^{-\frac{1}{2}(\mathbf{p}-\boldsymbol{\mu}^{\prime})^{\top}\mathbf{{{\Sigma}^{\prime}}^{-1}}(\mathbf{p}-\boldsymbol{\mu}^{\prime})}.$ (4)
@@
Figure 2: The underlying principles in handling mirrored contents. a) Mirror-NeRF employs a technique involving ray tracing and sampling to achieve mirror reflection. b) 3DGS mistakenly considers objects reflected by the mirror to be placed behind it, resulting the floaters behind the mirror. c) By employing depth and mirror mask loss on 3DGS, Evoluted-3DGS constrains the floaters directly onto the mirror surface. d) Our proposed Mirror-3DGS, captures the reflected objects by simulating a novel viewpoint situated behind the mirror.
3.2 Mirror-aware 3D Gaussian Representation
In the real world, the color of a point varies when observed from different viewpoints due to its intrinsic material and the environmental lighting. Although NeRF and 3DGS respectively utilize view-dependent MLPs and higher-order spherical harmonics to simulate this characteristic, the extreme specularity of mirrors amplifies such view-dependent inconsistency. Following, we review existing or potential solutions and introduce our reflection-agnostic Mirror-3DGS.
@@ 3.2 Mirror-aware 3D Gaussian Representation
$\bullet$ Evoluted-3DGS: We add mirror mask and depth constraints to the original 3DGS to avoid generating points behind mirrors, as depicted in Fig. 2(c). However, the limited mirror Gaussian spheres struggle to represent abundant scene reflection content, restricting the rendering quality of novel views.
$\bullet$ Mirror-3DGS (Ours): We innovatively introduce a mirror imaging method – “moving behind the mirror for a view of the world”, where the virtual image presented on the mirror is equivalent to observing the real world from behind the mirror, as illustrated in Fig. 2(d). Therefore, we avoid the time-consuming ray tracing required by Mirror-NeRF and also circumvent the insufficient expressiveness of Evoluted-3DGS’s mirror Gaussian spheres.
More specifically, 1) we render an image from the original viewpoint, i.e. observing from in front of the mirror, and retain the parts outside the mirror; 2) we mirror the original viewpoint to behind the mirror, rendering an image from the virtual viewpoint, and only keep the parts of the mirror (see Sec. 3.3); 3) finally, we merge these two images to form the final result (see Sec. 3.3).
@@
where $\mathbf{M}\in\mathbb{R}^{H\times W\times 3}$ represents the rendered mirror mask, and the rest of the symbols are defined as in Eq. (4). We utilize the GT mirror mask $\mathbf{M}_{gt}$ provided by the dataset to supervise the learning of the mirror attribute:
@@
3.3 Virtual Mirrored Viewpoint Construction and Image Fusion
Mirror Parameterization. In the previous section, we acquired the mirror attribute for each Gaussian. Then, based on the mirror attribute and opacity, we filter out the Gaussians belonging to the mirror for constructing a plane in the 3D space. We adopt the parameterization of an infinite plane [16]: $\pi=(\boldsymbol{n}_{\pi}^{\top},d)$ , where $\boldsymbol{n}_{\pi}=(a,b,c)^{\top}$ represents the normal of the plane, and $\frac{d}{\left\|\boldsymbol{n}_{\pi}\right\|}$ represents the distance from the origin of the plane to the world coordinate origin. Thereby, any point $\boldsymbol{p}=(x,y,z)^{\top}$ on the plane satisfies:
@@ 3.3 Virtual Mirrored Viewpoint Construction and Image Fusion
The optimization of the plane $\pi$ is presented in Sec. 3.4.
@@
The derivation of this matrix is detailed in the supplementary material.
Image Fusion. The image from the mirror viewpoint is obtained using the same rendering equation. Thus, we acquire the rendered image $\mathbf{C}_{\mathbf{P_{o}}}\in\mathbb{R}^{H\times W\times 3}$ from the current original viewpoint, the rendered image $\mathbf{C}_{\mathbf{P_{m}}}\in\mathbb{R}^{H\times W\times 3}$ from the mirror virtual viewpoint, and the mirror mask $\mathbf{M}$ for the current viewpoint. The final composite image are fused as follows:
@@
3.4 Two-Stage Training Strategy
We employ a two-stage training strategy, as illustrated in Fig. 3 .
Figure 3: The illustration of our two-stage training pipeline. In the Stage 1, supervision involves the mirror mask and the view image where mirror content is replaced with red color, focusing on learning the mirror plane and a coarse 3D Gaussian representation of the scene. In the Stage 2, based on the estimated mirror plane equation, views observed from the original viewpoint and its mirrored viewpoint are fused to form the final rendered result, further optimizing 3D Gaussians of the scene.
Stage 1: The training objective of the first stage is to obtain a rough Gaussian representation without mirror content and to optimize the mirror equation. We use the GT mirror mask provided by the dataset to erase the mirror content in the GT image $\mathbf{C}_{gt}$ and fill it with red, denoted as $\mathbf{C}_{wo}^{mirror}$. This is done to optimize the mirror attribute of Gaussians and to avoid generating false-Gaussians inside the mirror. The loss for rendering RGB images is consistent with 3DGS[18], i.e., L1 loss and D-SSIM term:
$\mathcal{L}_{rgb}^{1}=(1-\gamma)\mathcal{L}_{1}(\mathbf{C}_{\mathbf{P_{o}}},\mathbf{C}_{wo}^{mirror})+\gamma\mathcal{L}_{D-SSIM}(\mathbf{C}_{\mathbf{P_{o}}},\mathbf{C}_{wo}^{mirror}),$ (11)
where $\gamma$ is the weight, and $\mathbf{C}_{\mathbf{P_{o}}}$ is the rendered image from the original viewpoint, as defined in Sec. 3.3.
Additionally, to improve the quality of rendering and geometric structure, we add depth supervision. First, a pretrained depth estimation model [9] customized for mirror scenes is utilized to extract the depth map $\mathbf{D}_{gt}$ from image $\mathbf{C}_{wo}^{mirror}$. Then a modified rasterization library [34] is used to render the predicted depth map $\mathbf{D}_{pred}$, supervised with a similar loss:
$\mathcal{L}_{depth}=(1-\gamma)\mathcal{L}_{1}(\mathbf{D}_{pred},\mathbf{D}_{gt})+\gamma\mathcal{L}_{D-SSIM}(\mathbf{D}_{pred},\mathbf{D}_{gt}).$ (12)
The core of the first stage is to optimize the mirror equation $\pi=(\boldsymbol{n}_{\pi}^{\top},d)$ to ensure accurate generation of virtual mirrored viewpoints. We filter the mirror attribute by thresholding to obtain Gaussians with high mirror probability, and remove noise Gaussians through the improved RANSAC algorithm [11]. Then we supervise that the remaining Gaussians fall on the mirror plane:

### v2

@@ Abstract
3D Gaussian Splatting (3DGS) has significantly advanced 3D scene reconstruction and novel view synthesis. However, like Neural Radiance Fields (NeRF), 3DGS struggles with accurately modeling physical reflections, particularly in mirrors, leading to incorrect reconstructions and inconsistent reflective properties. To address this challenge, we introduce Mirror-3DGS, a novel framework designed to accurately handle mirror geometries and reflections, thereby generating realistic mirror reflections. By incorporating mirror attributes into 3DGS and leveraging plane mirror imaging principles, Mirror-3DGS simulates a mirrored viewpoint from behind the mirror, enhancing the realism of scene renderings. Extensive evaluations on both synthetic and real-world scenes demonstrate that our method can render novel views with improved fidelity in real-time, surpassing the state-of-the-art Mirror-NeRF, especially in mirror regions.
Index Terms:
Gaussian Splatting, Mirror Scene.
I Introduction
@@
footnotetext: means equal contribution and means corresponding authors. This work was supported in part by the National Natural Science Foundation of China no. 61931014, U21B2012, in part by the Beijing Natural Science Foundation no. L242014, and in part by the Fundamental Research Funds for the Central Universities.
@@
Traditional methods [1, 2, 3, 4, 5] simulate surface reflections by decomposing material properties and lighting conditions to handle scenes with mirrors. However, they often struggle to accurately infer surfaces and mirror components, especially in managing pure mirror reflections. With advancements in Neural Radiance Fields (NeRF), new methods [2, 6, 7] have emerged. NeRFReN [6] enhances rendering quality by using independent radiance fields for reflective and transmissive parts. Mirror-NeRF [7] employs Whitted Ray Tracing [8] for physically accurate mirror reflections, achieving high realism in novel view synthesis. Despite these improvements, NeRF-based methods face challenges such as long training times and slow rendering speeds, limiting their real-time application.
@@
Figure 1: The synthesized novel views and depth maps. 3DGS confuses mirror reflection, resulting in erroneous mirror depth. Ours properly handles content in the mirror with depth map almost identical to GT.
@@
Recently, 3D Gaussian Splatting (3DGS) [9] has emerged as a promising approach for real-time, high-fidelity 3D scene rendering. Instead of the slow volumetric rendering of NeRF, 3DGS explicitly represents scenes using 3D Gaussians and rasterizes 3D Gaussians onto the image plane using splatting techniques, achieving significant improvements in training and rendering speeds [10]. However, 3DGS tends to treat mirror reflections as separate scenes, constructing Gaussian points for them as if they physically exist, leading to inaccurate capture of these virtual images within mirrors during reconstruction and rendering, affecting the realism and visual coherence of the scene, as shown in Fig. 1.
@@
To address these challenges, we propose Mirror-3DGS, a novel rendering framework based on 3DGS, aimed at producing high-fidelity novel views in mirror-containing scenes. Our Mirror-3DGS is centered on the principle that the virtual image presented on the mirror is equivalent to observing the real world from behind the mirror. In general, We first identify 3D Gaussians representing the mirror, then observe from both the current and mirrored viewpoints, integrating these perspectives to render the final image. We introduce an attribute to signify mirror-like properties to 3D Gaussians and estimate the mirror plane equation, allowing us to derive the mirror transformation matrix and mirrored camera parameters. Our method involves a two-stage training process: 1) During the first stage, We learn a rough 3D Gaussian representation of the scene without mirror content and estimate the mirror plane equation. The mirror content is replaced with a fixed color based on the ground truth mirror mask to better learn mirror attributes and avoid interference from virtual images. 2) In the second stage, We filter Gaussians with high mirror properties for precise estimation of the mirror equation. By integrating views from both the current and mirrored viewpoints, we optimize the rendering quality of the entire scene.To ensure stability and improve the accuracy of plane estimation, we introduce planar consistency and depth constraints throughout the training process. These constraints help achieve more realistic and refined rendering results. Experiments on three synthetic and three real-world scenes demonstrate that Mirror-3DGS delivers rendering quality comparable to the state-of-the-art Mirror-NeRF, while significantly reducing training time and enabling real-time rendering. Our contributions can be summarized as follows:
@@
  • Innovative Rendering Framework for Mirror Scenes. We present Mirror-3DGS, a novel rendering framework that addresses the limitations of 3DGS in handling mirror scenes, enabling realistic synthesis of novel views.
@@
  • Extensive Experiments. We conduct comprehensive experiments across synthetic and real scenes to validate our proposed method. The results demonstrate our superiority in real-time rendering of high-quality novel views, particularly surpassing the state-of-the-art in mirror regions.
II Methodology
II-A Preliminaries
3D Gaussian Splatting (3DGS) is a state-of-the-art approach in novel view synthesis, utilizing the splatting technique [11] for efficient real-time rendering. In 3DGS, the scene is represented using anisotropic Gaussians characterized by center positions $\boldsymbol{\mu}\in\mathbb{R}^{3}$, covariances $\mathbf{\Sigma}\in\mathbb{R}^{3\times 3}$, color defined by spherical harmonic (SH) coefficients $\boldsymbol{c}\in\mathbb{R}^{3\times(k+1)^{2}}$ (with $k$ denoting the SH order), and opacity $\alpha\in\mathbb{R}^{1}$. The 3D Gaussian can be queried as follows:
$G(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\boldsymbol{\mu})^{\top}\mathbf{\Sigma}^{-1}(\mathbf{x}-\boldsymbol{\mu})},$ (1)
@@ II-A Preliminaries
$\boldsymbol{\mu}^{\prime}={\mathbf{P}\mathbf{W}\boldsymbol{\mu}},\quad\mathbf{{\Sigma}^{\prime}}=\mathbf{J}\mathbf{W}\mathbf{\Sigma}\mathbf{W^{\top}}\mathbf{J^{\top}},$ (2)
where $\boldsymbol{\mu}^{\prime}$ and $\mathbf{{\Sigma}^{\prime}}$ separately represent the 2D mean position and covariance of the projected 3D Gaussian. $\mathbf{P}$, $\mathbf{W}$ and $\mathbf{J}$ denote the projective transformation, viewing transformation, and Jacobian of the affine approximation of $\mathbf{P}$, respectively. The color of the pixel on the image plane, denoted by $\mathbf{p}$, uses a typical neural point-based rendering [13, 14]. Let $\mathbf{C}\in\mathbb{R}^{H\times W\times 3}$ represent the color of rendered image where $H$ and $W$ represents the height and width of images, the rendering process outlined as follows:
$\begin{split}\sigma_{i}=\alpha_{i}e^{-\frac{1}{2}(\mathbf{p}-\boldsymbol{\mu}^{\prime})^{\top}\mathbf{{{\Sigma}^{\prime}}^{-1}}(\mathbf{p}-\boldsymbol{\mu}^{\prime})}.\\ {\mathbf{C}(\mathbf{p})}={\sum_{i=1}^{N}c_{i}\sigma_{i}\prod_{j=1}^{i-1}(1-\sigma_{j})},\end{split}$ (3)
@@
Figure 2: The underlying principles in handling mirrored contents. a) Mirror-NeRF employs a technique involving ray tracing and sampling to achieve mirror reflection. b) 3DGS mistakenly considers objects reflected by the mirror to be placed behind it, resulting the floaters behind the mirror. c) Our proposed Mirror-3DGS, captures the reflected objects by simulating a novel viewpoint situated behind the mirror.
II-B Mirror-aware 3D Gaussian Representation
In the real world, a point’s color varies with different viewpoints due to its material and environmental lighting. While NeRF and 3DGS use view-dependent MLPs and spherical harmonics to simulate this, mirrors’ extreme specularity amplifies view-dependent inconsistencies. We review existing solutions and introduce our reflection-agnostic Mirror-3DGS.
@@ II-B Mirror-aware 3D Gaussian Representation
$\bullet$ Mirror-3DGS (Ours): We innovatively introduce a mirror imaging method – “moving behind the mirror for a view of the world”, where the virtual image presented on the mirror is equivalent to observing the real world from behind the mirror, as illustrated in Fig. 2(c). Therefore, we avoid the time-consuming ray tracing required by Mirror-NeRF.
More specifically, 1) we render an image from the original viewpoint, i.e. observing from in front of the mirror, and retain the parts outside the mirror; 2) we mirror the original viewpoint to behind the mirror, rendering an image from the virtual viewpoint, and only keep the parts of the mirror (see Sec. II-C); 3) finally, we merge these two images to form the final result (see Sec. II-C).
@@
where $\mathbf{M}\in\mathbb{R}^{H\times W\times 1}$ represents the rendered mirror mask, and the rest of the symbols are defined as in Eq. (3). We utilize the GT mirror mask $\mathbf{M}_{gt}$ provided by the dataset to supervise the learning of the mirror attribute:
@@
II-C Virtual Mirrored Viewpoint Construction and Image Fusion
Mirror Parameterization. In the previous section, we acquire the mirror attribute for each Gaussian. Using this attribute and opacity, we filter out Gaussians belonging to the mirror to construct a plane in 3D space. We adopt the parameterization of an infinite plane [15]: $\pi=(\boldsymbol{n}_{\pi}^{\top},d)$ , where $\boldsymbol{n}_{\pi}=(a,b,c)^{\top}$ represents the normal of the plane, and $\frac{d}{\left\|\boldsymbol{n}_{\pi}\right\|}$ represents the distance from the origin of the plane to the world coordinate origin. Thereby, any point $\boldsymbol{p}=(x,y,z)^{\top}$ on the plane satisfies:
@@ II-C Virtual Mirrored Viewpoint Construction and Image Fusion
The optimization of the plane $\pi$ is presented in Sec. II-D.
@@
Image Fusion. The image from the mirror viewpoint is obtained using the same rendering equation. We acquire the rendered image $\mathbf{C}_{\mathbf{P_{o}}}\in\mathbb{R}^{H\times W\times 3}$ from the original viewpoint, the rendered image $\mathbf{C}_{\mathbf{P_{m}}}\in\mathbb{R}^{H\times W\times 3}$ from the mirror viewpoint, and the mirror mask $\mathbf{M}$. The final composite image is fused as follows:
@@
II-D Two-Stage Training Strategy
We use a two-stage training strategy, as illustrated in Fig. 3.
Figure 3: The illustration of our two-stage training pipeline. In the Stage 1, we use the mirror mask and view image with mirror content replaced by red to learn the mirror plane and a coarse 3D Gaussian representation of the scene. In the Stage 2,based on the estimated mirror plane, we fuse views from the original and mirrored viewpoints to form the final rendered result, further optimizing the 3D Gaussians.
Stage 1: The training objective of the first stage is to obtain a rough Gaussian representation without mirror content and to optimize the mirror equation. We use the GT mirror mask provided by the dataset to erase the mirror content in the GT image $\mathbf{C}_{gt}$ and fill it with red, denoted as $\mathbf{C}_{wo}^{m}$. This is done to optimize the mirror attribute of Gaussians and to avoid generating false-Gaussians inside the mirror. The loss for rendering RGB images is consistent with 3DGS[9], i.e., L1 loss and D-SSIM term:
$\mathcal{L}_{rgb}^{1}=(1\!-\!\gamma)\mathcal{L}_{1}(\mathbf{C}_{\mathbf{P_{o}}},\mathbf{C}_{wo}^{m})+\gamma\mathcal{L}_{D-SSIM}(\mathbf{C}_{\mathbf{P_{o}}},\mathbf{C}_{wo}^{m})$ (10)
where $\gamma$ is the weight, and $\mathbf{C}_{\mathbf{P_{o}}}$ is the rendered image from the original viewpoint, as defined in Sec. II-C.
Additionally, to improve the quality of rendering and geometric structure, we add depth supervision. First, a pretrained depth estimation model [16] customized for mirror scenes is utilized to extract the depth map $\mathbf{D}_{gt}$ from image $\mathbf{C}_{wo}^{m}$. Then a modified rasterization library [17] is used to render the predicted depth map $\mathbf{D}_{pred}$, supervised with a similar loss:
$\mathcal{L}_{depth}=\mathcal{L}_{1}(\mathbf{D}_{pred},\mathbf{D}_{gt}).$ (11)
The first stage focuses on optimizing the mirror equation $\pi=(\boldsymbol{n}_{\pi}^{\top},d)$ to accurately generate virtual mirrored viewpoints. We filter the Gaussians with high mirror probability and use an improved RANSAC algorithm [18] to remove noise Gaussians. Then we supervise that the remaining Gaussians fall on the mirror plane: